If the [OH^ - ] of a solution is 2.46 * 10 ^ - 2 * M , what is the pH of the solution ?


1
Expert's answer
2021-05-03T07:42:28-0400

Concentration of "[OH^{-}]=2.46\times10^{-2}\space M"

"p[OH^-]=-log(2.46\times10^{-2})=1.60"

"p[H]+p[OH^-]=14"

"p[H]=14-1.60=12.4"